# The taggedPBC: Annotating a massive parallel corpus for crosslinguistic investigations

### v1

@@ Abstract
Existing datasets available for crosslinguistic investigations have tended to focus on large amounts of data for a small group of languages or a small amount of data for a large number of languages. This means that claims based on these datasets are limited in what they reveal about universal properties of the human language faculty. While this has begun to change through the efforts of projects seeking to develop tagged corpora for a large number of languages, such efforts are still constrained by limits on resources. The current paper reports on a large automatically tagged parallel dataset which has been developed to partially address this issue. The taggedPBC contains more than 1,800 sentences of pos-tagged parallel text data from over 1,500 languages, representing 133 language families and 111 isolates, dwarfing previously available resources. The accuracy of tags in this dataset is shown to correlate well with both existing SOTA taggers for high-resource languages (SpaCy, Trankit) as well as hand-tagged corpora (Universal Dependencies Treebanks). Additionally, a novel measure derived from this dataset, the N1 ratio, correlates with expert determinations of word order in three typological databases (WALS, Grambank, Autotyp) such that a Gaussian Naive Bayes classifier trained on this feature can accurately identify basic word order for languages not in those databases. While much work is still needed to expand and develop this dataset, the taggedPBC is an important step to enable corpus-based crosslinguistic investigations, and is made available for research and collaboration via GitHub.
@@ 1 Introduction
There are over 7,000 languages in the world, and such languages are rather diverse in terms of their typological features. At the same time, despite the overall diversity there are many similarities between languages, suggesting the possibility of universal properties that guide language evolution. These diversities and similarities have been identified and categorized in various ways by a variety of databases, most notably WALS (Dryer & Haspelmath 2013), [^1] https://wals.info/ Grambank (Skirgård et al., 2023), [^2] https://grambank.clld.org/ Autotyp (Bickel et al., 2023), [^3] https://github.com/autotyp/autotyp-data Glottolog (Hammarström et al. 2024), [^4] https://www.glottolog.org and Lexibank (List et al. 2022). [^5] https://lexibank.clld.org/ There are also collections of corpora that are expansive in their coverage of individual or groups of languages yet focused on particular genres or domains (CHILDES, [^6] https://childes.talkbank.org/ Brown, etc; and see the extensive corpora available via the CLARIN [^7] https://www.clarin.eu/resource-families project). The former tend to provide information about languages, while the latter tend to provide actual language data in various forms. The issue with the former is often lack of coverage for particular features for all languages in the dataset, whereas the latter typically suffer from insufficient annotation (many, for example, are monolingual or bilingual only) and lack linguistic coverage. Both kinds of datasets have a particular kind of value, but do not go far enough to address the concerns of linguistic diversity (i.e. Benítez-Burraco 2025) and gradience (i.e. Levshina et al. 2023) needed to expand our understanding of (cross-)linguistic processes.
@@
Recent advances in computational tools suggest the possibility of using such tools for automatic annotation, greatly reducing the time and effort needed to develop multilingual tagged datasets. For example, there is some indication that automated taggers for under-resourced languages can benefit from high-resource language part of speech tags in parallel corpora (Agic et al., 2015; Imani et al., 2022b), particularly for major word classes like nouns and verbs, even in zero-shot situations on low-resource languages. But while the use of such tools may greatly reduce the annotation burden for researchers, the derived dataset must also be validated for crosslinguistic investigation. Specifically, it is incumbent on researchers to identify a set of criteria used to evaluate whether the automatic annotations are sufficiently aligned with human-curated annotations or classifications before such a dataset can be considered robust enough for further crosslinguistic investigation.
@@
The current paper reports on the development of a large corpus of parallel texts tagged for word class that partly addresses the issues identified above. With data sourced from the Parallel Bible Corpus (PBC; Mayer & Cysouw 2014), computational word alignment tools allow for automatic tagging of word classes in 1,597 languages (roughly 20% of the world’s languages) that represent 133 families and 111 isolates.141414‘Isolates’ is defined here as languages found in the corpus that do not share family membership with another language in the corpus. The automated part of speech tagging method outlined in this paper was validated by comparing the results with output from existing automatic taggers for high-resource languages (SpaCy, Trankit) [^15] https://spacy.io/ ; http://nlp.uoregon.edu/trankit as well as with gold-standard tagged data in the UDT. The results indicate good alignment with both kinds of evaluation, providing a baseline from which to conduct crosslinguistic investigations. Further, features extracted from this dataset align with word order classifications in typological databases.
@@
Statistical properties of the derived corpus also comport with expectations regarding normality of language data. Importantly, this is partly because each language also has a large amount of data available for comparison. Individual corpora in the majority of these languages (over 1,500) contain more than 1,800 parallel sentences/verses, and the smallest corpus contains more than 700 sentences/verses. This dwarfs both the linguistic data and the language diversity available from other sources and the parallel nature of the data means that crosslinguistic investigations using this dataset can be confident that they are comparing similar contexts. The dataset (the taggedPBC) as well as explanations regarding its development and Python code for replicating the results described here, can be found at the linked repository. [^16] https://github.com/lingdoc/taggedPBC
@@
In the sections below I first outline the methodology used to develop the auto-taggers for each language (section 2). I then highlight statistics of the dataset (section 3) and describe tests used to assess tag validity (section 3.1). I also show that this dataset allows for extraction of features associated with typological properties, identifying a novel measure derived from this dataset (the “N1 ratio”) that correlates with word order classifications in multiple typological databases (section 3.2). I conclude the paper with discussion of ways the dataset could be improved to support more robust crosslinguistic investigations (section 4.1) as well as potential avenues of research made possible by this dataset (section 4.2).
@@ 2.1 Improving tag accuracy
There are several additional steps that I take to improve tagging accuracy, which fall under two general principles. The first principle is related to ensuring that parallel verses share as many common concepts as possible, i.e. that they are in some sense more semantically accessible. Increasing the number of shared concepts between languages in a dataset relative to rare concepts should allow a statistical word alignment model to more successfully identify alignments between shared concepts and/or words. This is akin to “under-sampling” in machine learning, whereby relevant properties of a dataset are adjusted to reduce noise in training. The second principle is related to ensuring that language-specific properties of languages (with regards to word formation) are taken into account. Since languages differ in terms of fusional properties (how many morphemes or units of meaning a word contains), this ought to be considered by any attempt at word alignment.
Accordingly, to address the concerns present under the first principle, I reduce the number of verses in the corpus using the following criteria. I first lemmatize two English New Testament translations, then reduce the dataset to verses which share at least 4 lemmas in those translations. This ensures that verses are more likely to be semantically accessible. Translations chosen for this comparison are the New Living Translation (NLT) and the New International Version (NIV), after discussions with Bible translators who work with low-resource languages. Translators indicated that these two versions are often consulted for their readability and accessibility in terms of how they translate the original Greek text, particularly in backtranslation (a checking step before final translation verification). The smaller dataset is then further reduced to verses where a verb occurs that is also present in at least five other verses, which increases the likelihood that each individual non-English Bible translation has common concepts that can be aligned with words in the English text. This procedure results in 1,885 verses for training word alignment models. As noted above, this is a principled method for under-sampling the parallel data, which should help improve the quality of the trained statistical word alignment model.
@@
It should also be noted that the languages in the PBC use several different writing systems and scripts. Some, like Devanagari, are syllabic, while others, like Arabic, are largely phonetic, and yet others, like Chinese, are systems that use individual characters to represent words or syllables. In order to ensure that we can adequately compare languages in terms of their sound structure, and to support the training of subword tokenizers, I convert all disparate systems to romanized script using the uroman [^17] https://www.isi.edu/~ulf/uroman.html tool, with the additional step of performing word-level tokenization for translations using scripts where words are not clearly separated by whitespace (fugashi [^18] https://github.com/polm/fugashi for Japanese, khmernltk [^19] https://github.com/VietHoang1512/khmer-nltk for Khmer/Kuy, jieba [^20] https://github.com/fxsjy/jieba for Chinese, pyidaungsu [^21] https://github.com/kaunghtetsan275/pyidaungsu for Burmese/Karen/Mon, attacut [^22] https://github.com/PyThaiNLP/attacut for Thai, botok [^23] https://github.com/OpenPecha/Botok for Tibetan) before converting to romanized script.
As a final step, I train each model with the source being the non-English language and the target being English. This enables a direct mapping of the words in the source language verse to the English words, allowing us to extract POS tags easily from words in the POS-tagged English sentence (via the SpaCy library). [^24] For the purposes of cross-linguistic comparison, I employ the Universal Dependencies tag set for parts of speech which includes 6 open classes (ADJ ‘adjective’, ADV ‘adverb’, INTJ ‘interjection’, NOUN ‘noun’, PROPN ‘proper noun’, VERB ‘verb’), 8 closed classes (ADP ‘adposition’, AUX ‘auxiliary’, CCONJ ‘coordinating conjunction’, DET ‘determiner’, NUM ‘numeral’, PART ‘particle’, PRON ‘pronoun’, SCONJ ‘subordinating conjunction’) and 3 ‘other’ classes for handling characters (PUNCT ‘punctuation’, SYM ‘symbol’, X ‘other’). These 17 classes, while not used by all languages, provide enough of a basic tag set to allow for cross-linguistic comparison. This methodology enables taggging of 1885 verses for a large number of languages in the PBC. Due to some differences in how sentences/clauses are translated between languages, and some portions not being available for some languages, there are a few mismatches. We can examine the actual number of sentences and sort them into bins, which gives counts of verses/languages in the corpus (Table 1).
@@ 2.2 Caveats
While automated transfer of POS tags from high-resource languages such as English to low-resource languages has been shown to give decent results, as noted by the literature above, there are some necessary caveats. First, limiting the number of sentences does remove some signal along with noise, meaning that it is possible not all word types will be represented in a given language. Additionally, mapping infrequent word types or function words, which are often less prominent and more functionally diverse crosslinguistically, will likely be less succcessful. At the same time, the large number of parallel sentences for the vast majority of languages in this dataset, and the common method by which they were developed, means that findings are statistically robust. As I demonstrate below, despite the caveats presented here, information and features extracted from the dataset comport with expected observations and lead to novel findings. This dataset is therefore released with the intention of providing a baseline for future research, with potential for expansion and refinement.
@@ 3 Results
The final taggedPBC dataset contains tagged data for 1,597 modern languages (1,599 if you include the 2 conlangs with NT translations: Esperanto and Klingon). This represents roughly 20% of the world’s languages from 133 distinct families and 111 isolates (i.e. languages without other family members in the dataset). The number of languages with more than 1,800 sentences/verses in the dataset is 1,545 (see Table X), and the fewest number of sentences/verses is represented by Tanimuca-Retuarã/Letuama (ISO 639-3: tnc) with 754 sentences. The automatic tagging method outlined above results in a statistically robust number of unique arguments and predicates for all 1597 languages (mean: 732 arguments, 495 predicates), with the fewest number of unique arguments and predicates being tagged for Cantonese (ISO 639-3: yue; 163 arguments, 74 predicates).
As noted above, this dwarfs all available crosslinguistic tagged data in terms of language coverage, with the closest similar corpus-based dataset being the UDT, in which only 106 of the 160 languages in the v2.14 release have more than 200 tagged sentences. However, before we can use this dataset for investigation, we need to be confident in the validity of the tags as well as the features that can be extracted from this dataset. Accordingly, we now turn to validation of the tagging accuracy of the IBM2 word alignment POS-tag transfer method (section 3.1), as well as how well extracted features correlate with existing classifications in typological databases (section 3.2).
@@ 3.1 Validating tag accuracy
There are two means by which we can validate the automated tags provided by the IBM2 word alignment method. The first involves assessing against other automatic taggers developed for high resource languages. Here we want to see that the method used for tagging these languages aligns with existing language-specific taggers at a level above chance. The second means of assessment involves comparing the words and their tags with hand-tagged data in gold standard corpora. Here, if we find exact matches (in form and tag) between the taggedPBC corpora and those in hand-tagged corpora, we can be confident that other words in the taggedPBC with the same part of speech do in fact belong to that word class.

### v2

@@ Abstract
Existing datasets available for crosslinguistic investigations have tended to focus on large amounts of data for a small group of languages or a small amount of data for a large number of languages. This means that claims based on these datasets are limited in what they reveal about universal properties of the human language faculty. While this has begun to change through the efforts of projects seeking to develop tagged corpora for a large number of languages, such efforts are still constrained by limits on resources. The current paper reports on a large tagged parallel dataset which has been developed to partially address this issue. The taggedPBC contains POS-tagged parallel text data from more than 1,940 languages, representing 155 language families and 78 isolates, dwarfing previously available resources. The accuracy of particular tags in this dataset is shown to correlate well with both existing SOTA taggers for high-resource languages (SpaCy, Trankit) as well as hand-tagged corpora (Universal Dependencies Treebanks). Additionally, a novel measure derived from this dataset, the N1 ratio, correlates with expert determinations of intransitive word order in three typological databases (WALS, Grambank, Autotyp) such that a Gaussian Naive Bayes classifier trained on this feature can accurately identify basic intransitive word order for languages not in those databases. While much work is still needed to expand and develop this dataset, the taggedPBC is an important step to enable corpus-based crosslinguistic investigations, and is made available for research and collaboration via GitHub.
@@ 1 Introduction
There are over 7,000 languages in the world, and such languages are rather diverse in terms of their typological features. At the same time, despite the overall diversity there are many similarities between languages, suggesting the possibility of universal properties that guide language evolution. These diversities and similarities have been identified and categorized in various ways by a variety of databases, most notably WALS (Dryer & Haspelmath 2013), [^1] https://wals.info/ Grambank (Skirgård et al., 2023), [^2] https://grambank.clld.org/ Autotyp (Bickel et al., 2023), [^3] https://github.com/autotyp/autotyp-data Glottolog (Hammarström et al. 2024), [^4] https://www.glottolog.org and Lexibank (List et al. 2022). [^5] https://lexibank.clld.org/ There are also collections of corpora that are expansive in their coverage of individual or groups of languages yet focused on particular genres or domains (CHILDES, [^6] https://childes.talkbank.org/ Brown, etc; and see the extensive corpora available via the CLARIN [^7] https://www.clarin.eu/resource-families project). The former tend to provide information about languages, while the latter tend to provide actual language data in various forms. The issue with the former is often lack of coverage for particular features for all languages in the dataset, whereas the latter typically suffer from insufficient annotation and lack linguistic coverage. Both kinds of datasets have a particular kind of value, but do not go far enough to address the concerns of linguistic diversity (i.e. Ben´ıtez-Burraco 2025) and gradience (i.e. Levshina 2023) needed to expand our understanding of (cross-)linguistic processes.
@@
Recent advances in computational tools suggest the possibility of using such tools for automatic annotation, greatly reducing the time and effort needed to develop multilingual tagged datasets. As an example, there is some indication that automated taggers for under-resourced languages can benefit from high-resource language part of speech tags in parallel corpora (Agic et al., 2015; Imani et al., 2022b), particularly for major word classes like nouns and verbs, even in zero-shot situations on low-resource languages. But while the use of such tools may greatly reduce the annotation burden for researchers, the derived dataset must also be validated for crosslinguistic investigation. Specifically, it is incumbent on researchers to identify a set of criteria used to evaluate whether the automatic annotations are sufficiently aligned with human-curated annotations or classifications before such a dataset can be considered robust enough for further crosslinguistic investigation.
@@
The current paper reports on the development of a large corpus of parallel texts tagged for word class that partly addresses the issues identified above. With data sourced from the Parallel Bible Corpus (PBC; Mayer & Cysouw 2014: 1,597 languages) and additional data sourced from newly-available Bible translations (345 languages), computational word alignment tools allow for automatic tagging of word classes in 1,942 languages (roughly 25% of the world’s languages) that represent 155 families and 78 isolates.141414‘Isolates’ is defined here as languages found in the corpus that do not share family membership with another language in the corpus. The automated part of speech tagging method outlined in this paper was validated by comparing the results with output from existing automatic taggers for high-resource languages (SpaCy, Trankit) [^15] https://spacy.io/ ; http://nlp.uoregon.edu/trankit as well as with gold-standard tagged data in the UDT. The results indicate good alignment with both kinds of evaluation, providing a baseline from which to conduct crosslinguistic investigations. Further, features extracted from this dataset align with word order classifications in typological databases.
@@
Statistical properties of the derived corpus also comport with expectations regarding normality of language data. Importantly, this is partly because each language also has a large amount of data available for comparison. Individual corpora in the majority of these languages (over 1,850) contain more than 1,800 parallel sentences/verses, and the smallest corpus contains more than 700 sentences/verses. This dwarfs both the linguistic data and the language diversity available from other sources and the parallel nature of the data means that crosslinguistic investigations using this dataset can be confident that they are comparing similar contexts. The dataset (the taggedPBC) as well as explanations regarding its development and Python code for replicating the results described here, can be found at the linked repository. [^16] https://github.com/lingdoc/taggedPBC
@@
In the sections below I first outline the methodology used to develop the auto-taggers for each language (§2). I then highlight statistics of the dataset (§3) and describe tests used to assess tag validity (§3.1). I also show that this dataset allows for extraction of features associated with typological properties, identifying a novel measure derived from this dataset (the “N1 ratio”) that correlates with intransitive word order classifications in multiple typological databases (§3.2). I conclude the paper with discussion of ways the dataset could be improved to support more robust crosslinguistic investigations (§4.1) as well as potential avenues of research made possible by this dataset (§4.2).
@@ 2.1 Improving tag accuracy
There are several additional steps that I take to improve tagging accuracy, which fall under two general principles. The first principle is related to ensuring that parallel verses share as many common concepts as possible, i.e. that they are in some sense more semantically accessible. Increasing the number of shared concepts between languages in a dataset relative to rare concepts should allow a statistical word alignment model to more successfully identify alignments between shared concepts and/or words. This is akin to “under-sampling” in machine learning, whereby relevant properties of a dataset are adjusted to reduce noise in training. The second principle is related to ensuring that specific properties of individual languages (with regards to word formation) are taken into account. Since languages differ in terms of fusional properties (how many morphemes or units of meaning a word contains), this ought to be considered by any attempt at word alignment.
Accordingly, to address the concerns present under the first principle, I reduce the number of verses in the corpus using the following criteria. I first lemmatize two English New Testament translations, then reduce the dataset to verses which share at least 4 lemmas in those translations. This ensures that verses are more likely to be semantically accessible. Translations chosen for this comparison are the New Living Translation (NLT) and the New International Version (NIV), after discussions with Bible translators who work with low-resource languages. Translators indicated that these two versions are often consulted for their readability and accessibility in terms of how they translate the original Greek text, particularly in backtranslation (a checking step before final translation verification). The smaller dataset is then further reduced to verses where a verb occurs that is also present in at least five other verses, which increases the likelihood that each individual non-English Bible translation has common concepts that can be aligned with words in the English text. This procedure results in 1,884 verses for training word alignment models. As noted above, this is a principled method for under-sampling the parallel data, which should help improve the quality of the trained statistical word alignment model.
@@
It should also be noted that the languages in the PBC use several different writing systems and scripts. Some, like Devanagari, are syllabic, while others, like Arabic, are largely phonetic, and yet others, like Chinese, are systems that use individual characters to represent words or syllables. In order to ensure that we can adequately compare languages in terms of their sound structure, and to support the training of subword tokenizers, I convert all disparate systems to romanized script using the uroman [^17] https://www.isi.edu/~ulf/uroman.html tool, with the additional step of performing word-level tokenization for translations using scripts where words are not clearly separated by whitespace (fugashi [^18] https://github.com/polm/fugashi for Japanese, khmernltk [^19] https://github.com/VietHoang1512/khmer-nltk for Khmer/Kuy, jieba [^20] https://github.com/fxsjy/jieba for Chinese, pyidaungsu [^21] https://github.com/kaunghtetsan275/pyidaungsu for Burmese/Karen/Mon, attacut [^22] https://github.com/PyThaiNLP/attacut for Thai, botok [^23] https://github.com/OpenPecha/Botok for Tibetan) before converting to romanized script. [^24] As noted by a reviewer, this is not an ideal mapping, since romanization does not necessarily reflect the phonemic values of a language in the same way that a standardized orthography does. Putting aside the question of phonemic or phonetic accuracy, however, I would argue that romanization does at least allow for comparison in a way that retaining original scripts does not.
As a final step, I train each model with the source being the non-English language and the target being English. This enables a direct mapping of the words in the source language verse to the English words, allowing us to extract POS tags easily from words in the POS-tagged English sentence (via the SpaCy library). That is, once the respective translation model has been trained for a particular language, every word of every sentence/verse is back-translated to English. If the translation of that word matches a word in the respective POS-tagged English sentence/verse, the source language word is tagged with the English POS. [^25] For the purposes of cross-linguistic comparison, I employ the Universal Dependencies tag set for parts of speech which includes 6 open classes (ADJ ‘adjective’, ADV ‘adverb’, INTJ ‘interjection’, NOUN ‘noun’, PROPN ‘proper noun’, VERB ‘verb’), 8 closed classes (ADP ‘adposition’, AUX ‘auxiliary’, CCONJ ‘coordinating conjunction’, DET ‘determiner’, NUM ‘numeral’, PART ‘particle’, PRON ‘pronoun’, SCONJ ‘subordinating conjunction’) and 3 ‘other’ classes for handling characters (PUNCT ‘punctuation’, SYM ‘symbol’, X ‘other’). These 17 classes, while not used by all languages, provide enough of a basic tag set to allow for cross-linguistic comparison. This methodology enables taggging of 1,884 verses for a large number of languages in the PBC. Due to some differences in how sentences/clauses are translated between languages, and some portions not being available for some languages, there are a few mismatches. We can examine the actual number of sentences and sort them into bins, which gives counts of verses/languages in the corpus (Table 1).
@@ 2.2 Caveats
While automated transfer of POS tags from high-resource languages such as English to low-resource languages has been shown to give decent results, as noted by the literature above, there are some necessary caveats. First, limiting the number of sentences does remove some signal along with noise, meaning that it is possible not all word types will be represented in a given language. Along with this, mapping infrequent word types or function words (i.e. that in English, der in German), which are often less prominent and more functionally diverse crosslinguistically, will likely be less succcessful. Another case in point is the presence of such terms as numeral classifiers: many languages require special words or markers that classify nouns being enumerated, but such a class of words is simply not present in other languages, like English. Since we are transferring tags from English, word classes like numeral classifiers will not be tagged in this dataset at all, which means there is scope for improving on these annotations. At the same time, the large number of parallel sentences for the vast majority of languages in this dataset, and the common method by which they were developed, means that findings for other word types (specifically nouns, proper nouns, pronouns, verbs, and auxiliaries) are statistically robust.
The methodology followed here should not be construed as a claim that all languages share the same tag set, but rather that certain word classes (nouns, verbs, etc.) exist in the majority of languages and can be readily identified via translation. At the same time, there are a large number of words in each of these corpora that simply have no tag (more than 50% of tokens in most corpora are tagged as “unk”) because there is no direct translation equivalent corresponding to a tagged English word in the same verse. This reflects the methodological goal of maximizing accuracy of POS transfer. While this leaves potentially ambiguous words for future investigation (and see the discussion in the concluding section below regarding development of language-specific automated tools or manual annotation for individual low-resource corpora), the translation method provides a decent proxy measure that supports the identification of major word classes. As I demonstrate below, despite the caveats presented here, information and features extracted from this tagged dataset comport with expected observations and lead to novel findings. This dataset is therefore released in order to provide a baseline for future research, with potential for expansion and refinement, particularly in relation to language-specific POS tags.
@@ 3 Results
The final taggedPBC dataset contains tagged data for 1,942 modern languages (1,944 if you include the 2 conlangs with NT translations: Esperanto and Klingon). This represents roughly 25% of the world’s languages from 155 distinct families and 78 single languages or isolates. The number of languages with more than 1,800 sentences/verses in the dataset is 1,853 (see Table 1), and the fewest number of verses is represented by Nafusi (ISO 639-3: jbn) with 729 verses. The automatic tagging method outlined above results in a statistically robust number of unique arguments (combining nouns, proper nouns, pronouns) and predicates (combining verbs and auxiliaries) for all 1,942 languages (mean: 699 arguments, 466 predicates), with the fewest number of unique arguments being tagged for Nugunu (ISO 639-3: yas; 76) and the fewest number of unique predicates being tagged for Rapanui (ISO 639-3: rap; 57). If we compare the number of unique nouns and verbs, the fewest unique nouns are identified for ut-Hun (ISO 639-3: uth; 59) and the fewest unique verbs for Bambalang (ISO 639-3: bmo; 54).
As noted above, this is significantly more than currently available crosslinguistic tagged data in terms of language coverage, with the closest similar corpus-based dataset being the UDT, in which only 106 of the 160 languages in the v2.14 release have more than 200 tagged sentences. However, before we can use this dataset for investigation, we need to be confident in the validity of the tags as well as the features that can be extracted from this dataset. Accordingly, we now turn to validation of the tagging accuracy of the IBM2 word alignment POS-tag transfer method (§3.1), as well as how well extracted features correlate with existing classifications in typological databases (§3.2).
@@ 3.1 Validating tag accuracy
There are two means by which we can validate the automated tags provided by the IBM2 word alignment method. The first involves assessing against other automatic taggers developed for high resource languages. Here we want to see that the method used for tagging these languages aligns with existing language-specific taggers at a level above chance. The second means of assessment involves comparing the words and their tags with hand-tagged data in gold standard corpora. Here, if we find exact matches (in both word form and tag) between the taggedPBC corpora and those in hand-tagged corpora, we can be confident that other words in the taggedPBC with the same part of speech do in fact belong to that word class. For the following evaluations I focus strictly on the word classes of nouns and verbs, as these are most well-represented in the data.